OpenStudy (anonymous):

Solve for m. m 9 = 25 m A) ± 216 B) ± 225 C) ± 14.6 D) ± 15

OpenStudy (anonymous):

This is a proportion. It should look like m/9 = 25/m

Parth (parthkohli):

1) Multiply both sides by 9. 2) Multiply both sides by m. 3) Solve the quadratic equation that you get.

OpenStudy (lgbasallote):

cross multiply \[\frac m9 = \frac{25}{m}\] \[m^2 = 25\times 9\] \[m^2 = 225\] \[m = \pm \sqrt{225}\] \[m = \pm 15

OpenStudy (theviper):

Cross multiply :- \[\Large{\frac{m}{9}=\frac{25}{m}}\]\[\Large{m \times m=25 \times 9}\]\[\LARGE{{m^2=225}}\]

OpenStudy (anonymous):

what do you mean by both sides?

Parth (parthkohli):

Never mind, that is not a quadratic equation.

OpenStudy (lgbasallote):

\[m = \pm 15\] is what i meant in my last line

OpenStudy (theviper):

\[\Huge{m=\pm 15}\]

OpenStudy (anonymous):

OHHHHHH OKAY! I get it lol I miss-read it

OpenStudy (anonymous):

cross multiplication method is easy for you.

OpenStudy (theviper):

Understood??

OpenStudy (anonymous):

Thanks @muhammad9t5 @TheViper @Igbasallote and @ParthKohli !!!!!!

Parth (parthkohli):

You're welcome! :)

OpenStudy (theviper):

Most Welcome :)

OpenStudy (anonymous):

I would not actually do the multiplication of 9 times 25… FWIW